OpenStudy (anonymous):

How would you write 5x^20-7x^10+15 in quadratic form?

jimthompson5910 (jim_thompson5910):

hint: let z = x^10. Squaring both sides gives z^2 = x^20

jimthompson5910 (jim_thompson5910):

(x^10)^2 = x^(10*2) = x^20

OpenStudy (anonymous):

Hmm would it be -5(x^10)² - 7(x^10) + 15?

jimthompson5910 (jim_thompson5910):

then replace every x^10 with z

jimthompson5910 (jim_thompson5910):

the first 5 is positive and not negative

OpenStudy (anonymous):

Can anyone help I'm stuck between the answers A and C.

jimthompson5910 (jim_thompson5910):

you didn't list the answer choices